OpenStudy (salina):

Ananda wanted to determine the average of her 6 test scores.She added the scores correctly to get T ,but divided by 7 instead of 6 .The result was 12 less than her actual average . Which equation coule be use to determine the value of T?

OpenStudy (anonymous):

T/6 - T/7 = 12

OpenStudy (anonymous):

what you do is this. We basically have two equations and two unknowns. Let x be the average that she incorrectly calculated, and let T be, well, T. so we have: x+12 = T/6 and x=T/7 so T=7x, substitute into the first equation, solve for x, then solve for 7. or do it LA-MrC's way

OpenStudy (anonymous):

Multiply the entire equation by 42. 7T - 6T = 504 T = 504

OpenStudy (anonymous):

The erroneous average was 72. The correct average is 84.

OpenStudy (salina):

don't solve just write equation only